OpenStudy (anonymous):

use properties of logarithms to find the exact value of the expression. 9^(log(base9)1-log(base9)2) type an integer or a simplified fraction.

OpenStudy (anonymous):

I think: \[x^{\log_x(a)} = a\] I am not sure whether I am true or not..

OpenStudy (anonymous):

If above one is right then: We will have: \[\log(a) - \log(b) = \log(\frac{a}{b})\] use this firstly..

OpenStudy (anonymous):

Here a = 1 and b = 2... Can you use this formula??

OpenStudy (anonymous):

so it will be \[9^{\log \frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

well i mean log base 9

OpenStudy (anonymous):

Yep but don't forget to write base 9.. So it will be: \[\huge 9^{\log_9(\frac{1}{2})}\] Okay.. Now use the first formula that I gave.

OpenStudy (anonymous):

It is similar to \(x^{log_{x}(a)}\)..

OpenStudy (anonymous):

\[9^{\log _{9}\frac{ 1 }{ 2 }}=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Yep, according to me this should be the answer...

OpenStudy (anonymous):

so it should be 1/2??

OpenStudy (anonymous):

or the entire equation?

OpenStudy (anonymous):

You have reduced entire expression to 1/2. That is it..!!

OpenStudy (anonymous):

oh! okay thank you!!!!!

OpenStudy (anonymous):

You are welcome dear...